# Region-Aware Text-to-Image Generation via Hard Binding and Soft Refinement

### v1

@@ 4.3 Ablation Study
We introduced a parameter $r$ to control the times of hard binding. As shown in Figure 7, excessive hard binding reduces the times for subsequent soft refinement, weakening its impact. This may diminish interactions between adjacent objects, potentially causing noticeable boundaries to reappear. Our tests indicate that setting $r$ between 1 and 3 typically achieves an ideal balance, effectively combining positional control with seamless integration between regions.
Effectiveness of Factor $\delta$.
$\delta$ controls the proportion of the original prompt. As shown in Figure 7, when $\delta=0$, indicating no soft refinement, the results exhibit noticeable partitioning. As the weight adjusted by soft refinement increases, image fusion improves gradually. In our extensive testing, $\delta$ proves to be a flexible parameter, with values in the range (0,1] consistently producing high-quality images. The optimal $\delta$ can be adjusted according to specific scenes and generation requirements.
@@
Since RAG employs decoupled multi-region generation operations, each region is processed independently during the generation process. This results in an increase in inference time as the number of regions grows, as shown in Figure 8. However, because RAG does not alter the original model’s architecture, existing acceleration techniques, such as hyper-flux [24] , can still be applied to RAG. By incorporating these acceleration methods, inference time can be significantly reduced.Moreover, after the introduction of these acceleration techniques, the region generation process in RAG remains unaffected, ensuring that the precision and detail control for each region remain stable. This demonstrates that, although inference speed presents a challenge, with appropriate acceleration strategies, RAG can still efficiently and accurately generate multi-region images while maintaining both generation quality and control over the interactions between regions.
@@ 5 Conclusion
In this paper, we introduce RAG, a novel training-free region-aware text-to-image generation method designed to address challenges posed by regional prompts. RAG operates through two key stages. First, regional hard binding independently constructs the content of each region. Then, regional soft refinement enhances interactions between adjacent regions and improves attribute generation. Furthermore, RAG bypasses the need for external inpainting models, enabling direct image repainting to modify unsatisfactory regions from last generations. Extensive experimental results demonstrate the superiority of RAG on compositional generation compared to prior training-free methods.

### v2

@@ 4.3 Ablation Study
Figure 7: Qualitative analysis of hard binding steps $r$ and soft refinement strength $\delta$. A few steps of binding is sufficient for regional completeness, and appropriate soft refinement intensity leads to improved regional coherence.
Figure 8: RAG, accelerated by Hyper-Flux, is 2.5$\times$ faster than the original version, maintaining visual quality.
@@
We introduced a parameter $r$ to control the frequency of hard binding applications. As shown in Figure 7, excessive hard binding limits the opportunity for subsequent soft refinement, weakening its impact. This may diminish interactions between adjacent objects, potentially causing noticeable boundaries to reappear. By adjusting $r$, we aim to balance precise object placement with smooth integration, optimizing both the structure and cohesion of the generated image.
Effectiveness of Coefficient $\delta$.
We introduced the coefficient $\delta$ to modulate the intensity of regional soft refinement. As illustrated in Figure 7, excessively low or high $\delta$ values may result in noticeable partitioning or slight misalignment between text and image. Setting $\delta$ to an optimal level enhances the fusion of image regions, yielding a more coherent and natural overall output.
@@
In RAG, the decoupled multi-region generation operation processes each region independently, leading to increased inference time as the number of regions grows. However, because RAG does not modify the underlying architecture of the original model, existing acceleration frameworks, such as Hyper-Flux [25], remain compatible with RAG. As shown in Figure 8, combining RAG with Hyper-Flux results in inference being 2.5-fold faster while preserving the stability of accuracy and detail control in regional generation. This demonstrates that, although inference speed presents a challenge, with appropriate acceleration strategies, RAG can efficiently generate multi-region images with high quality and maintain control over interactions between regions.
@@ 5 Conclusion
In this paper, we introduce RAG, a novel tuning-free Region-Aware text-to-image Generation method designed to address challenges posed by regional prompts. RAG operates through two key stages. First, regional hard binding independently constructs the content of each region. Then, regional soft refinement enhances interactions between adjacent regions and improves attribute generation. Furthermore, RAG bypasses the need for external inpainting models, enabling direct image repainting to modify unsatisfactory regions from last generations. Extensive experimental results demonstrate the superiority of RAG on compositional generation compared to prior tuning-free methods.